Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Please help solve

Answer by checkley75(3666)

3c+g=175
2c+4g=200 now multipy the first equation by -4 & add thus
-12c-4g=-700
-------------------------------------
-10c=-500
c=-500/-10
c=50 cost of 1 chocolate bar. now substitute 50 for c & solve for g.
3*50+g=175
150+g=175
g=175-150
g=25 cost of 1 pack of gum.
proof
2*50+4*25=200
100+100=200
200=200